# RainPro-8: An Efficient Deep Learning Model to Estimate Rainfall Probabilities Over 8 Hours

### v1

@@ 5 Conclusion & Future Work
We introduce RainPro-8, a deep learning model for precipitation nowcasting in Europe that outperforms existing operational systems and deep learning nowcasting models. RainPro-8 addresses the challenges of forecasting sparse and stochastic precipitation by probabilistic predictions at different precipitation levels. It leverages multiple data sources with an efficient model that forecasts all lead times simultaneously, while considering the ordinality of precipitation levels.
Going forward, RainPro-8 could be extended to be more robust to structural changes in input data and to missing data — the latter is of particular importance in operational settings where some input source could become absent. RainPro-8 uses GFS forecasts as part of its inputs, where it would be interesting to identify alternative data sources, and study automatically learned thresholds.
@@ A.2 Data Preprocessing
Each data source is modified to match the desired resolution for model input, either through upsampling, such as GFS data from 28km/px to 16km/px, or downsampling, like radar data from 2km/px to both 4km/px and 8km/px. The time dimension also presents challenges to consider. Satellite data, for instance, has a 1-hour operational delay, making the most recent timestep available at -60 minutes. For the source gfs_16km, we select 122 channels representing multiple weather variables at different pressure levels (Appendix F). As GFS is only initialized 4 times a day, we use the latest forecast lead time as our initial state. The subsequent lead times form the gfs_forecast_16km, from which we use only the precipitation rate variable. Table 3 summarizes all data sources with the corresponding spatial and temporal details.
@@
[…]
@@ B.2 Additional Evaluation
Table 6: Frequency Bias Index (FBI) across different thresholds for precipitation forecasting models. Our model, RainPro-8, has a tendency to over-predict high-precipitation events, in contrast to GFS, NowcastNet, Earthformer, and SimVP, which consistently under-predict rainfall. PySTEPS appears the most balanced in terms of bias, with over- and under-predictions occurring at similar rates—though this does not translate to higher forecast accuracy.
@@ Appendix D Visualizations
Figure 14: Sample ground truth, RainPro-8, MetNet-3*, GFS, and PySTEPS, NowcastNet, Earthformer, and SimVP forecasts at selected lead times with origin at 2024-01-23 11:20 UTC.
@@
Figure 16: Sample ground truth, RainPro-8, MetNet-3*, GFS, and PySTEPS, NowcastNet, Earthformer, and SimVP forecasts at selected lead times with origin at 2024-04-19 06:00 UTC.
@@
Figure 18: Sample ground truth, RainPro-8, MetNet-3*, GFS, and PySTEPS, NowcastNet, Earthformer, and SimVP forecasts at selected lead times with origin at 2024-06-28 11:20 UTC.
@@ Appendix E Intensity Classes
Appendix F GFS Input Variables
@@
Appendix G Compute Resources
Training is performed on an NVIDIA H100 80GB SXM5 GPU with 26 vCPUs and 225 GiB RAM. Table 9 shows the training time for each experiment including baselines and ablation studies.
Table 9: .
Experiment Training Time (h)
RainPro-8 13:36
MetNet-3* 18:16
NowcastNet 6:48
Earthformer 24:03
SimVP 6:36
RainPro-8_ce 15:52
RainPro-8_clt 18:06
RainPro-8_noltw 13:21
RainPro-8_nomask 13:20
RainPro-8_sources 13:22

### v2

@@ 5 Conclusion & Future Work
We introduce RainPro-8, a deep learning model for precipitation nowcasting that outperforms existing operational systems and deep learning nowcasting models. RainPro-8 addresses the challenges of forecasting probabilistic predictions at different precipitation levels. It leverages multiple data sources with an efficient model that forecasts all lead times simultaneously, exploiting ordinality of precipitation levels. Going forward, RainPro-8 could be extended to explicitly address robustness to missing data which is of particular importance in operational settings where some input source could become absent, and to automatically learn thresholds.
6 Reproducibility Statement
The code will be made publicly available, along with instructions for accessing data via the official sources. The paper provides comprehensive details on the RainPro-8 framework, architecture, and evaluation, while the appendices serve as supporting material. Specifically, Section 3 describes the task definition, loss function, output generation, and model architecture. Section 4 summarizes the training setup and hyperparameters (with further details in Appendices E and F), data sources (including preprocessing and sample details in Appendices A and I), and baselines, metric definitions, and performance comparisons (expanded in Appendices B and G). Section 4 also presents visualizations (also in Appendix D), ablation studies, attribution analyses (also in Appendix C), and an evaluation of a RainPro-8 variant adapted for the benchmark task on a standard precipitation nowcasting benchmark (also in Appendix H).
@@ A.2 Data Preprocessing
Each data source is modified to match the desired resolution for model input, either through upsampling, such as GFS data from 28km/px to 16km/px, or downsampling, like radar data from 2km/px to both 4km/px and 8km/px. The time dimension also presents challenges to consider. Satellite data, for instance, has a 1-hour operational delay, making the most recent timestep available at -60 minutes. For the source gfs_16km, we select 122 channels representing multiple weather variables at different pressure levels (Appendix I). As GFS is only initialized 4 times a day, we use the latest forecast lead time as our initial state. The subsequent lead times form the gfs_forecast_16km, from which we use only the precipitation rate variable. Table 4 summarizes all data sources with the corresponding spatial and temporal details.
@@
We tackle challenges such as varying data ranges and skewed distribution of precipitation intensities using normalization, clipping, and binning. Min-max normalization is applied to all data sources, as many variables are non-normally distributed. Radar reflectivity values (dBZ) are clipped to a range of -1 to 64 to remove outliers. To address rainfall skewness, target radar maps are categorized into predefined rain intensity classes for multi-label classification. Using the Marshall-Palmer equation (Marshall & Palmer, 1948), dBZ values are converted to mm/h, with finer granularity for light rain ($\geq 0.1,\geq 0.2,\geq 0.4$) and broader ranges for heavier rainfall ($\textrm{up to}\geq 25.0\ \textrm{mm/h}$) (Appendix E). This approach captures subtle light rain variations while accounting for rare intense precipitation events, aligning with the original dBZ distribution.
@@ B.2 Additional Evaluation
Table 7: Frequency Bias Index (FBI) across different thresholds for precipitation forecasting models. Our model, RainPro-8, has a tendency to over-predict high-precipitation events, in contrast to GFS, Earthformer, and SimVP, which consistently under-predict rainfall. PySTEPS appears the most balanced in terms of bias, with over- and under-predictions occurring at similar rates—though this does not translate to higher forecast accuracy.
@@ Appendix D Visualizations
Figure 15: Sample ground truth, RainPro-8, MetNet-3*, GFS, and PySTEPS, RainPro-8R, Earthformer, and SimVP forecasts at selected lead times with origin at 2024-01-23 11:20 UTC.
@@
Figure 17: Sample ground truth, RainPro-8, MetNet-3*, GFS, and PySTEPS, RainPro-8R, Earthformer, and SimVP forecasts at selected lead times with origin at 2024-04-19 06:00 UTC.
@@
Figure 19: Sample ground truth, RainPro-8, MetNet-3*, GFS, and PySTEPS, RainPro-8R, Earthformer, and SimVP forecasts at selected lead times with origin at 2024-06-28 11:20 UTC.
@@ Appendix E Intensity Classes
Appendix F Compute Resources
Training is performed on an NVIDIA H100 80GB SXM5 GPU with 26 vCPUs and 225 GiB RAM. Table 9 shows the training and inference times for each experiment, including baselines and ablation studies.
Table 9: Training and inference times for all models trained.
Experiment Training Time (h) Inference Time (s/sample)
RainPro-8 13:36 0.03
MetNet-3* 18:16 2.12
RainPro-8R 13:22 0.03
Earthformer 24:03 0.03
SimVP 6:36 0.12
RainPro-8 w/ CE loss (no OC) 15:52 0.03
RainPro-8 w/ lead time conditioning 18:06 2.10
RainPro-8 w/o lead time weights 13:21 0.03
Appendix G Robustness Analysis
The primary experiments in this work were conducted with a single training run due to the substantial computational resources required. While this setup provides a fair basis for comparison, the closeness of certain results motivates an analysis of the variability that may arise from different random seeds. Therefore, we conducted three independent runs for the two most comparable experiments: RainPro-8 and its cross-entropy variant, RainPro-8 w/ CE loss (no OC). Table 10 reports the mean and standard deviation across these runs for multiple evaluation metrics. The results show that both models converge consistently, with RainPro-8 demonstrating a slight but consistent advantage across all metrics.
Table 10: Aggregated results from three independent runs (mean ± std). Best results are highlighted in bold.
Metric RainPro-8 RainPro-8 w/ CE loss (no OC)
CRPS 0.06096 ± 0.00004 0.06099 ± 0.00002
CSI 0.27923 ± 0.00054 0.27832 ± 0.00032
FSS 0.53663 ± 0.00119 0.53532 ± 0.00041
FBI 1.26048 ± 0.01224 1.26999 ± 0.00242
MAE 0.12593 ± 0.00102 0.12610 ± 0.00043
MSE 1.50463 ± 0.01819 1.52142 ± 0.01898
Appendix H SEVIR Nowcasting Experiments
H.1 Dataset and Prediction Task
SEVIR (Veillette et al., 2020) is a widely used benchmark for precipitation nowcasting, containing 20,393 weather events of radar frame sequences of 4-hour duration covering $384\times 384~\textrm{km}^{2}$ (1 km/pixel) at 5-minute intervals. Following Diffcast (Yu et al., 2024), we formulate the prediction task as forecasting 20 frames given 5 initial frames (5 $\rightarrow$ 20), and the spatial resolution is downsampled to $128\times 128$ pixels to reduce computational cost.
H.2 RainPro-2R
We modify RainPro-8 to create RainPro-2R, a radar-only variant suitable for short-term nowcasting. RainPro-2R uses only radar frames as input, without additional data sources or spatial context with respect to the targets. The model predicts 20 frames at a downsampled resolution of $128\times 128$ pixels, while retaining the same three downsampling layers used in RainPro-8. Training follows the Diffcast (Yu et al., 2024) pipeline, employing our ordinal consistent loss and single-step prediction approach. RainPro-2R is trained using the Adam optimizer with a learning rate of 0.0001 for a total of 200K iterations, and the model with the lowest validation loss is selected for evaluation.
The intensity classes that define the probability distribution are based on the SEVIR benchmark thresholds: [0.0, 16.0, 31.0, 59.0, 74.0, 100.0, 133.0, 160.0, 181.0, 219.0, 255.0]. Given the shorter lead times, the decay rate for the lead time weights is set to 2.
H.3 Evaluation
To assess the accuracy of precipitation nowcasting, we evaluate RainPro-2R against both deterministic and stochastic generative models. Deterministic baselines include SimVP (Gao et al., 2022) and Earthformer (Gao et al., 2024b), which employ a recurrent-free strategy to generate all frames simultaneously, as well as PhyDNet (Guen & Thome, 2020), MAU (Chang et al., 2021), and ConvGRU (Shi et al., 2017a), which generate frames sequentially using recurrent strategies. For state-of-the-art generative baselines, we include STRPM (Chang et al., 2022), MCVD(Voleti et al., 2022), and PreDiff (Gao et al., 2024a).
Evaluation metrics follow prior work (Yu et al., 2024) and include Critical Success Index (CSI), Heidke Skill Score (HSS), pooled CSI, and visual perception metrics LPIPS Zhang et al. (2018) and SSIM Wang et al. (2004). In addition, we introduce the previously discussed Fraction Skill Score (FSS) and Continuous Ranked Probability Score (CRPS).
Details on CSI, CRPS, and FSS are in Appendix B.1. Pooled CSI consists of computing CSI after applying max-pooling to the predinctions and targets. Therefore, pooled CSI only checks whether any pixel within a neighborhood is hit, ignoring how many pixels match.
The models are evaluated at thresholds of [16, 74, 133, 160, 181, 219] (Yu et al., 2024). Pooled CSI is computed over 4×4 and 16×16 neighborhoods, and CRPS over 5×5 and 17×17 neighborhoods, to have 2 or 8 pixels in each direction.
Heidke Skill Score (HSS)
HSS measures forecast accuracy while accounting for correct hits expected by chance. The values range between -0.5 and 1.0, with scores greater than 0 indicating skill by improving over a random forecast.
$\text{HSS}=\frac{2\,(\text{TP}\cdot\text{TN}-\text{FN}\cdot\text{FP})}{(\text{TP}+\text{FN})(\text{FN}+\text{TN})+(\text{TP}+\text{FP})(\text{FP}+\text{TN})},$ (12)
where TP, FP, and FN are true positives, false positives, and false negatives, based on the threshold (rate $\geq$ threshold).
Learned Perceptual Image Patch Similarity (LPIPS)
LPIPS (Zhang et al., 2018) measures perceptual similarity between predicted and observed images using deep network features. It can capture differences in texture, structure, and high-level perceptual qualities. Lower LPIPS values indicate predictions that are perceptually closer to the observations.
Structural Similarity Index Measure (SSIM)
SSIM (Wang et al., 2004) evaluates image similarity by comparing luminance, contrast, and structural information between predicted and observed images. It reflects how well the predicted patterns match the true spatial structures. Higher SSIM values indicate better preservation of structural and visual information in the forecast.
Appendix I GFS Input Variables
@@
Appendix J LLM Usage
Large Language Models have been employed solely for proofreading and to polish writing.